A firm’s demand curve in period 1 is Q=25 - P. Fixed costs are 20 and marginal costs per unit are 5.



1
Expert's answer
2021-05-13T17:28:48-0400

Total revenue is Price×Quantity, Rearraging the demand equation, we get

"P=25-Q"


So,

Total Revenue"=PQ=(25-Q)\u00d7Q=25Q-Q^2"


Marginal Revenue would be differentiation of Total revenue. So,


Marginal revenue"=25-2Q"